# online_appendix

### v1

@@ Abstract
This paper documents a new feature of international trade: the increase in the number of low-value transactions. Using Spanish data, we show that the share of low-value transactions in the total number of transactions increased from 9% to 61% in exports and from 14% to 54% in imports between 1997 and 2023. The increase in the number of low-value trade transactions is related to the rise in e-commerce and direct-to-customer sales facilitated by online retail platforms. In the case of exports, the increase in the number of low-value transactions is also explained by the fast-fashion strategy followed by clothing firms.
@@ 1 Introduction
The new stage of the globalization process that began in the 1980s presents some characteristics, such as the revolution in information and communication technology, the emergence of China as a global manufacturing exporter, and the expansion of global value chains, which were unknown in previous integration episodes (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). This paper documents an additional novel feature of this globalization stage: the rise in the number of low-value transactions.
@@
Despite the large increase in the number of low-value transactions, they represent a very small share of the total value of exports and imports: 0.4% and 0.2%, respectively, in 2023. We show that clothing and footwear dominate low-value transactions in both exports and imports, and have contributed notably to the increase in the share of low-value transactions in the total number of trade transactions between 1997 and 2023. We document that countries geographically close to Spain are the main destinations, by volume, for low-value export transactions, while China is the most important origin, by volume, for Spanish low-value import transactions. Furthermore, China was the largest contributor to the rise in the number of low-value import transactions between 1997 and 2023. We also discover that there are fewer firms performing low-value transactions than large-value transactions. Finally, we find that the concentration of number of transactions by firm is much higher for low-value transactions than for high-value transactions. For example, the firm that performed the largest number of low-value transactions in 2023 represented 20% and 21% of the total number of low-value export and import transactions, respectively. Regarding high-value transactions, the top firm “only” represented 6% and 2% of the total number of high-value export and import transactions, respectively.
@@
Second, the increase in the number of low-value import transactions is related to the emergence of online retail platforms, such as Shein or Temu, which allow consumers to purchase a wide range of products while providing foreign manufacturers with access to a global market. Transactions mediated through these platforms are denoted as direct to consumer, because purchases go directly from the overseas producer to the local consumer (Fajgelbaum and Khandelwal,, 2024). Consistent with this explanation, we show that the increase in the number of low-value import transactions rocketed between 2021 and 2023, when the above-mentioned platforms became popular in Spain. Furthermore, these import transactions originated in China, where these online platforms were founded. The geographical distribution and administrative justification for the large number of low-value import transactions from China are also consistent with a direct-to-consumer shipment explanation. For earlier periods, we provide evidence consistent with the argument that online stores and marketplaces explain the increase in the number of low-value import transactions. Finally, we show that the rise in the number of low-value export transactions occurring from 2021 onward, which is not related to fast-fashion retailers, is also consistent with the emergence and development of online marketplaces.
@@
To the best of our knowledge, this paper is the first to document the increase in the share of low-value transactions in the total number of international trade transactions. This discovery adds to the literature that has analyzed the features of the globalization stage that began in the 1980s (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). We also add to the literature on the role of the de minimis threshold in international trade. Fajgelbaum and Khandelwal, (2024) show that the value of de minimis imports in the US increased from 0.05 billion USD in 2012 to 54.5 billion USD in 2023. In volume, US de minimis imports increased from 0.1 to 1 billion shipments between 2012 and 2023. They also show that it is the less affluent consumers who are more likely to be the recipients of low-value imports. We contribute to this literature by analyzing both low-value export and import transactions, quantifying the share in value and volume of low-value transactions in total trade transactions, exploring a longer time period, and documenting stylized facts about the number of low-value transactions, such as product composition, destinations and origins, number of traders, and concentration of transactions across traders.
@@
Our paper also speaks to the literature on e-commerce, online platforms, and direct-to-consumer shipments. Online platforms reduce barriers to finding and evaluating potential foreign consumers or suppliers, and facilitate logistics and international payments (Lendle et al.,, 2016; Chen and Wu,, 2021; Carballo et al.,, 2022). Volpe-Martincus et al., (2024) argue that the expansion of direct-to-consumer shipments is related to the increased participation of consumers in international trade. We contribute to this literature by providing evidence consistent with the argument that corporate online stores and online marketplaces contributed to the rise in the number of low-value trade transactions.
@@
Finally, we also contribute to the literature exploring the relationship between fast fashion and international trade. Fernandes and Tang, (2020) show that Portuguese clothing manufacturers responded to Chinese competition by increasing the frequency of their shipments. We show that the high frequency of shipments by fashion retailers played a very important role in the rise in the number of low-value export transactions. Furthermore, we show that these transactions were concentrated in countries that were geographically far from Spain and that the products were transported by air. These latter findings also link our paper to the literature that analyzes how the time sensitivity of demand determines the choice of transport means (Hummels and Schaur,, 2013) and the lumpiness of trade (Alessandria et al.,, 2010; Kropf and Sauré,, 2014; Hornok and Koren, 2015a, ; Hornok and Koren, 2015b, ; Békés et al.,, 2017).
@@ 2 Data and stylized facts
We obtain trade data on goods for the period 1997-2023 from the Customs and Excise Department of the Spanish Tax Agency (AEAT-Customs). Each observation included in the data set captures the commercial exchange of a product between a Spanish trade operator (a firm or an individual) and a foreign trade operator. Each observation provides information on the direction of the trade flow (exports or imports), the Spanish trade operator’s customs identification code, the product in the 8-digit Combined Nomenclature (CN) classification, the value (in euros), the country of destination or origin, the Spanish province of destination or origin, the month in which the transaction occurred, and the date the transaction was documented in customs. Our data set does not provide an identification of the foreign partner. We remove all transactions that do not involve the transfer of property in exchange for financial compensation. Hence, our dataset does not include transactions related to gifts, donations, or return of goods. [^2] We exclude Andorra and Gibraltar from the list of origins and destinations because, due to their location, trade with these territories is similar to domestic sales for Spanish firms.
We define a transaction as the exchange recorded by an observation included in our dataset: the commercial exchange of a product between a Spanish and a foreign trade operator. It is important to note that our dataset captures transactions and not shipments. Therefore, an import transaction of 150 euros or less from a non-EU country recorded in our database cannot be automatically equated with a de minimis shipment since this transaction could belong to a shipment that includes multiple products.
@@ C. Top 20 HS 2-digit chapters by share in total number of low-value transactions (2023)
The first stylized fact, whose documentation constitutes the main contribution of this paper, is the large increase in the share of low-value transactions in the total number of international trade transactions. It rises from 9% to 61% in exports and from 14% to 54% in imports. That is, in 2023, low-value transactions represented more than half of the total number of trade transactions. [^4] The median values of an export and import transaction in 1997 were 4,678 and 2,776 euros, respectively. In 2023, they decreased to 77 and 114 euros. As shown in Figure A.1 in the appendix, this key stylized fact is robust to using a low-value threshold adjusted for inflation, a low-value threshold based on the 800 USD de minimis value applied by the United States, and to splitting the sample into EU and non-EU members. [^5] The exception is that the share of low-value export transactions in the total number of transactions between 2021 and 2023 decreased in non-EU27 countries and increased in EU27 countries.
@@
The total value of low-value exports increased from 16 million euros in 1997 to 1,478 million euros in 2023 (multiplied by 92). For imports, the value of low-value transactions increased from 30 million euros to 735 million euros (multiplied by 24). These increases were much larger than those in total value, which multiplied by 4 for both exports and imports. Despite these large increases, our second stylized fact is that low-value transactions still represent a very small share of the value of trade. Panel B shows that low-value export transactions represented 0.4% and 0.2% of total export and import value in 2023, respectively. In 1997, the figures were 0.02% and 0.03% for exports and imports, respectively.
@@
A major limitation of our dataset is that most low-value transactions do not report the traded product. Specifically, 70% and 71% of low-value trade transactions in 2023 did not report the exported and imported product, respectively. This occurs because export transactions equal to or below the de minimis value of the destination country, and import transactions equal to or below 150 euros are not obliged to report the traded product. Keeping this limitation in mind, the third stylized fact is that clothing and footwear represented the largest share of low-value transactions that recorded the traded product in 2023: 41% and 30% of the total number of low-value export and import transactions, respectively (panel C of Figure 1). Electrical and electronic products, such as semiconductors or switchers, and computer parts also accounted for a sizable share of the total number of export and import transactions. By broad economic activities, consumption goods represented 67% and 53% of the total number of low-value export and import transactions, respectively, in 2023; intermediate goods represented 26% and 37% of export and import transactions, respectively; and parts related to capital goods represented 7% and 10% of export and import transactions, respectively. Interestingly, in 1997, books was the top product in exports and electrical machinery in imports (panel A of Figure A.2 in the appendix). The share of clothing and footwear in 1997 was 12% and 13% for exports and imports, respectively. This comparison highlights that the clothing and footwear industries played an important role in the growth of the number of low-value transactions. [^6] Panel A of Figure A.3 in the appendix shows that clothing and footwear also occupied prominent roles among the most exported and imported products in high-value transactions in 2023. However, their share in the total number of transactions was less than that in low-value transactions: 23% in large-value export transactions versus 41% in low-value transactions, and 22% versus 30% in import transactions.
@@ C. Concentration of the number of transactions by firm, 2023 (% of the total number of transactions)
China was the main source of low-value imports in 2023, accounting for 37% of the total number of import transactions. Only eight countries that appear on the list of the top 20 export destinations are also included in the list of the top 20 import origins. In particular, China, which is the most important origin of low-value imports by volume, is not included in the list of the top 20 destinations. [^7] If we added-up China and Hong-Kong, the ranking of the origins of low-value import transactions in 2021 would be similar to that in 2023 (Figure A.4 in the appendix). The ranking of export destinations is similar in 1997 and 2023 (panel B of Figure A.2 in the appendix). In contrast, China represented less than 5% of the total number of low-value import transactions in 1997. This comparison suggests that China played a key role in the growth of the number of low-value import transactions. [^8] Panel B of Figure A.3 shows that China also represented a large share of high-value import transactions in 2023. However, this share was smaller than that in low-value import transactions: 20% versus 37%. The increase in low-value imports from China has raised concerns in some countries. Policy makers have proposed to remove the de minimis threshold for Chinese products, arguing that it is used as a tool to avoid tariffs or the import of goods that do not meet health and safety standards. [^9] See “Where Textile Mills Thrived, Remnants Battle for Survival” by Jordyn Holman, The New York Times, 21 January 2024 at https://www.nytimes.com/2024/01/21/business/economy/textile-mills-carolina-trade-de-minimis.html and“EU takes aim at China’s Temu and Shein with proposed import duty” by Andy Bounds Paola Tamma, Financial Times, 3 July 2024 at https://www.ft.com/content/1c4c0bee-f67e-404b-877d-e0cb38faf2d6.
@@
We estimated descriptive regressions to identify the country-level factors correlated with having a high share in Spain’s total number of low-value export and import transactions. Table A.1 in the appendix shows that the number of low-value export transactions is positively correlated with GDP and GDP per capita of the importer, sharing a land border and official language, and negatively correlated with distance and being a member of the EU. The number of low-value import transactions is positively correlated with the GDP of the source country, being a member of the EU, and negatively correlated with sharing an official language. Belonging to a free trade agreement does not play any role in either exports or imports. This is consistent with the fact that low-value import transactions, by definition, are duty-free. [^10] Volpe-Martincus et al., (2024) get a similar result when comparing firms’ and consumers’ imports. The former would be equivalent to our high-value transactions and the latter to our low-value transactions. However, we observe that belonging to the EU has a large negative effect on the share of low-value export transactions in the total number of export transactions, while it has a large positive effect on imports.
@@
The fifth stylized fact, illustrated by panel B of Figure 2, is that the number of firms that only trade in low-value transactions is smaller than the number that trade in high-value transactions or in high- and low-value transactions. For example, in 2023, the number of firms that only traded in low-value transactions represented 8% and 16% of all exporters and importers, respectively. These percentages were much higher than those at the beginning of the period (2% for both exporters and importers). There is an increase in the percentage of exporters that trade in large- and low-value transactions. In contrast, the percentage of this category of firms decreases among importers. We find a positive correlation between the number of large- and low-value transactions performed by a firm both for exports and imports. We also document a positive correlation between the total value of large- and low-value transactions by a firm for both exports and imports (Table A.2 in the appendix).
@@
Finally, the sixth stylized fact, illustrated by panel C of Figure 2, is that the concentration of the number of transactions by firm is higher in low-value transactions than in high-value ones. For example, the top firm in number of low-value export transactions represented 20% of all low-value export transactions in 2023, while the top firm in high-value export transactions “only” represented 6% of total high-value export transactions. For imports, the differences were even larger: the top firm represented 21% and 2% of the total number of small- and high-value transactions, respectively. Furthermore, the concentration of transactions by firm increased between 1997 and 2023, in particular for low-value transactions: the share of the top firm in the total number of transactions increases from 3% to 20% in exports and from 3% to 21% in imports.

### v2

@@ Abstract
This paper documents a new feature of international trade: the increase in the share of low-value transactions in the total volume of transactions. Using Spanish data, we show that the share of low-value transactions in the total number of transactions increased from 9% to 61% in exports and from 14% to 54% in imports between 1997 and 2023. The increase in the number of low-value trade transactions is explained by the rise of e-commerce and direct-to-customer sales facilitated by online retail platforms, and the fast-fashion strategy followed by clothing firms.
@@ 1 Introduction
The new stage of the globalization process that began in the 1980s presents some characteristics, such as the revolution in information and communication technology, the emergence of China as a global manufacturing exporter, and the expansion of global value chains, which were unknown in previous integration episodes (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). This paper documents an additional novel feature of this globalization stage: the increase in the share of low-value transactions in the total number of international trade transactions.
@@
We show that clothing and footwear dominate low-value transactions in both exports and imports, and have contributed notably to the increase in the share of low-value transactions in the total number of trade transactions between 1997 and 2023. We document that China is the most important source for low-value import transactions and the largest contributor to the increase in the number of low-value import transactions between 1997 and 2023. We also discover that there are fewer firms performing low-value transactions than large-value transactions, and the concentration of the number of transactions by firm is much higher for low-value transactions than for high-value transactions.
Despite the large increase in the number of low-value transactions, they represent a small share of the total value of exports and imports: 0.4% and 0.2%, respectively, in 2023. However, if we used the less stringent de minimis threshold applied by the US, 800 USD, low-value transactions would represent a nonnegligible share of Spanish trade: 1.5% and 0.8% of the total value of exports and imports, and 3.8% and 2.6% of the value of exports and imports in consumer goods.
@@
Second, the increase in the number of low-value import transactions is related to the emergence of online retail platforms, such as Shein or Temu, that allow consumers to purchase a wide range of products while providing foreign manufacturers with access to a global market. Transactions mediated through these platforms are denoted as direct to consumer, because purchases go directly from the overseas producer to the local consumer (Fajgelbaum and Khandelwal,, 2024). Consistent with this explanation, we show that the increase in the number of low-value import transactions rocketed between 2021 and 2023, when the above-mentioned platforms became popular. Furthermore, these import transactions originated in China, where these online platforms were founded. The geographical distribution and administrative justification for the large number of low-value import transactions from China are also consistent with a direct-to-consumer shipment explanation. For earlier periods, we provide evidence consistent with the argument that online stores and marketplaces explain the increase in the number of low-value import transactions. Finally, we show that the increase in the number of low-value export transactions occurring from 2021 onward, which is not related to fast-fashion retailers, is also consistent with the emergence and development of online marketplaces.
Is the increase in the share of low-value transactions in the total number of trade transactions a particular feature of Spanish trade, or is it a global trend? Additional analyses seem to confirm the trend observed for Spain, at least for imports. Fajgelbaum and Khandelwal, (2024) show that the volume of de minimis imports in the United States multiplied by 10 between 2012 and 2013, and the share of de minimis imports in the total value of US imports increased from 0.002% to 1.720% during the same period. In the case of the EU, according to European Commission, (2023), the volume of e-commerce imports from countries outside the EU tripled between 2019 and 2022. Regarding exports, the fact that some Spanish firms are global leaders in the fashion industry could have enhanced the share of low-value transactions in the volume of export transactions.
The increase in the share of low-value transactions in the total number of transactions has important implications for international trade. First, a growing share of imports enter the US and the EU with a value equal to or below the de minimis threshold to avoid paying tariffs is argued. This concern is particularly strong in the US since tariffs on many Chinese goods were increased in 2018. Furthermore, there are claims that de minimis enables the import of goods that do not meet health and safety standards. [^2] See “Where Textile Mills Thrived, Remnants Battle for Survival” by Jordyn Holman, The New York Times, 21 January 2024 at https://www.nytimes.com/2024/01/21/business/economy/textile-mills-carolina-trade-de-minimis.html and “The EU targets China’s Temu and Shein with proposed import duty” by Andy Bounds Paola Tamma, Financial Times, 3 July 2024 at https://www.ft.com/content/1c4c0bee-f67e-404b-877d-e0cb38faf2d6. The alleged misuse of the de minimis threshold has led to proposals to eliminate it both in the United States and in the EU. Therefore, the increase in the share of low-value transactions could have important implications for trade policy.
Second, the increase in the share of low-value transactions has been facilitated by the emergence of e-commerce platforms that allow consumers to choose from a wide range of products. These platforms allow consumers to be directly involved in international trade, and, therefore, their purchasing decisions have a greater weight in shaping trade flows (Volpe-Martincus et al.,, 2024). If there are differences in the variables that govern consumer purchasing decisions from those of firms, it will be important to take those differences into account when designing economic policies. For example, Volpe-Martincus et al., (2024) show that low-value transactions are not affected by trade agreements and this paper shows that the time-sensitiveness of low-value transactions leads to a greater use of air transport.
Third, low-income consumers import more low-value products than high-income ones (Fajgelbaum and Khandelwal,, 2024). The increase in low-value import transactions leads to a greater direct participation of low-income consumers in trade and underscores the need to consider the impact of trade policies, such as the elimination of the de minimis threshold, on income inequality.
@@
To the best of our knowledge, this paper is the first to document the increase in the share of low-value transactions in the total number of international trade transactions. This discovery adds to the literature that has analyzed the features of the globalization stage that began in the 1980s (Krugman,, 1995; Rodrik,, 2011; Baldwin,, 2016; Antràs,, 2020). We also add to the literature on the role of the de minimis threshold in international trade. As mentioned above, Fajgelbaum and Khandelwal, (2024) document the rise in de minimis imports in the US and show that less affluent consumers are more likely to be recipients of low-value imports. We contribute to this literature by analyzing both low-value export and import transactions, quantifying the share in value and volume of low-value transactions in total trade transactions, exploring a longer time period, and highlighting some stylized facts about low-value transactions. We also offer a new explanation for the increase in the number of low-value transactions: the fast-fashion strategy followed by fashion retailers.
@@
Our paper also speaks to the literature on e-commerce, online platforms, and direct-to-consumer shipments. Online platforms reduce barriers to finding and evaluating potential foreign consumers or suppliers, and facilitate logistics and international payments (Lendle et al.,, 2016; Chen and Wu,, 2021; Carballo et al.,, 2022). We contribute to this literature by providing evidence consistent with the argument that corporate online stores and online marketplaces contributed to the rise in the number of low-value trade transactions.
@@
Finally, we also contribute to the literature exploring the relationship between fast fashion and international trade. Fernandes and Tang, (2020) show that Portuguese clothing manufacturers responded to Chinese competition by increasing the frequency of their shipments. We show that the high frequency of shipments by fashion retailers played a very important role in the increase in the number of low-value export transactions. Furthermore, we show that these transactions were concentrated in countries that were geographically far from Spain and products that were transported by air. These latter findings also link our paper to the literature that analyzes how the time sensitivity of demand determines the choice of transport means (Hummels and Schaur,, 2013) and the lumpiness of trade (Alessandria et al.,, 2010; Kropf and Sauré,, 2014; Hornok and Koren, 2015a, ; Hornok and Koren, 2015b, ; Békés et al.,, 2017).
@@ 2 Data and stylized facts
We obtain trade data on goods for the period 1997-2023 from the Customs and Excise Department of the Spanish Tax Agency (AEAT-Customs). Each observation captures the exchange of a product between a Spanish trade operator (a firm or an individual) and a foreign trade operator. It provides information on the direction of the trade flow (exports or imports), the Spanish trade operator’s customs identification code, the product in the 8-digit Combined Nomenclature (CN) classification, the value (in euros), the country of destination or origin, the Spanish province of destination or origin, the month in which the transaction occurred, and the date the transaction was documented in customs. It does not provide an identification of the foreign partner. We remove all transactions that do not involve the transfer of property in exchange for financial compensation. Hence, our dataset does not include transactions related to gifts, donations, or return of goods. [^3] We exclude Andorra and Gibraltar from the list of origins and destinations because, due to their location, trade with these territories is similar to domestic sales for Spanish firms.
We define a transaction as the exchange recorded by a single observation of our dataset: the commercial exchange of a product between a Spanish and a foreign trade operator. It is important to note that our dataset captures transactions and not shipments. Therefore, an import transaction of 150 euros or less from a non-EU country recorded in our database cannot be automatically equated with a de minimis shipment since this transaction could belong to a shipment that includes multiple products.
@@ C. Top 20 HS 2-digit chapters by share in total number of low-value transactions (2023)
The first stylized fact, whose documentation constitutes the main contribution of this paper, is the large increase in the share of low-value transactions in the total number of international trade transactions. It rises from 9% to 61% in exports and from 14% to 54% in imports. That is, in 2023, low-value transactions represented more than half of the total number of trade transactions. [^5] The median values of an export and import transaction in 1997 were 4,678 and 2,776 euros, respectively. In 2023, they decreased to 77 and 114 euros. As shown in Figure A.1 in the Appendix, this stylized fact is robust to using a low-value threshold adjusted for inflation, a low-value threshold based on the 800 USD de minimis value applied by the United States, and to splitting the sample into EU and non-EU members. [^6] The exception is that the share of low-value export transactions in the total number of transactions between 2021 and 2023 decreased in non-EU27 countries and increased in EU27 countries.
As mentioned above, in general, a transaction cannot be equated with a shipment. However, if some conditions are met, a transaction can be classified as a shipment. [^7] This happens when for a given export (or import) transaction, we cannot find another transaction that is performed by the same firm, happens in the same month, uses the same mode of transport, originates in the same province, is destined to the same country, and was recorded in customs on the same date. Although the percentage of transactions that can be classified as shipments is small, the evolution of the share of low-value shipments in total shipments between 1997 and 2023 is qualitatively similar to that found for transactions. [^8] The share of low-value shipments in total shipments multiplies by 3 both for exports and imports between 1997 and 2023.
@@
The total value of low-value exports increased from 16 million euros in 1997 to 1,478 million euros in 2023 (multiplied by 92). For imports, the value of low-value transactions increased from 30 million euros to 735 million euros (multiplied by 24). These increases were much larger than those in total value, which multiplied by 4 for both exports and imports between 1997 and 2023. Despite these large relative increases, our second stylized fact is that low-value transactions still represent a very small share of the total value of trade: 0.4% and 0.2% of the total value of exports and imports in 2023, respectively (panel B). [^9] In 1997, the figures were 0.02% and 0.03% for exports and imports, respectively. Fajgelbaum and Khandelwal, (2024) calculated that de minimis imports represented 7.2% of US imports of consumer goods in 2023. If we used the 800 USD threshold to define low-value transactions, these would represent 2.6% of the Spanish imports of consumer goods in 2023. For exports, the share of low-value transactions would increase to 3.8%. Therefore, when using the US de minimis threshold, low-value transactions represent a nonnegligible share of the Spanish trade in consumer goods. [^10] Using the US thresholds, low-value transactions would represent 1.5% and 0.8% of the total value of exports and imports in 2023.
@@
A major limitation of our dataset is that most low-value transactions do not report the traded product. Specifically, 70% and 71% of low-value trade transactions in 2023 did not report the exported or imported product, respectively. This occurs because export transactions equal to or below the de minimis value of the destination country, and import transactions equal to or below 150 euros are not obliged to report the traded product. Keeping this limitation in mind, the third stylized fact is that clothing and footwear represented the largest share of low-value transactions that recorded the traded product in 2023: 41% and 30% of the total number of low-value export and import transactions, respectively (panel C of Figure 1). By broad economic activities, consumption goods represented 67% and 53% of the total number of low-value export and import transactions, respectively, in 2023; intermediate goods represented 26% and 37% of export and import transactions, respectively; and parts related to capital goods represented 7% and 10% of export and import transactions, respectively. Interestingly, in 1997, books were the top product in exports and electrical machinery in imports (panel A of Figure A.2 in the appendix). The share of clothing and footwear in 1997 was 12% and 13% for exports and imports, respectively. This comparison highlights that the clothing and footwear industries played an important role in the growth of the number of low-value transactions. [^11] Panel A of Figure A.3 in the appendix shows that clothing and footwear also occupied prominent roles among the most exported and imported products in high-value transactions in 2023. However, their share in the total number of transactions was less than that in low-value transactions: 23% in large-value export transactions versus 41% in low-value transactions, and 22% versus 30% in import transactions.
@@ C. Concentration of the number of transactions by firm, 2023 (% of the total number of transactions)
China was the main source of low-value imports in 2023, accounting for 37% of the total number of import transactions. Only eight countries that appear on the list of the top 20 export destinations are also included in the list of the top 20 import origins. In particular, China, which is the most important origin of low-value imports by volume, is not included in the list of the top 20 destinations. [^12] If we added-up China and Hong-Kong, the ranking of the origins of low-value import transactions in 2021 would be similar to that in 2023 (Figure A.4 in the appendix). The ranking of export destinations is similar in 1997 and 2023 (panel B of Figure A.2 in the appendix). In contrast, China represented less than 5% of the total number of low-value import transactions in 1997. This comparison suggests that China played a key role in the growth of the number of low-value import transactions. [^13] Panel B of Figure A.3 shows that China also represented a large share of high-value import transactions in 2023. However, this share was smaller than that in low-value import transactions: 20% versus 37%.
@@
We estimated descriptive regressions to identify the country-level factors correlated with having a high share in Spain’s total number of low-value export and import transactions. Table A.1 in the appendix shows that the number of low-value export transactions is positively correlated with GDP and GDP per capita of the importer, sharing a land border and official language, and negatively correlated with distance and being a member of the EU. The number of low-value import transactions is positively correlated with the GDP of the source country, being a member of the EU, and negatively correlated with sharing an official language. Belonging to a free trade agreement does not play any role in either exports or imports. This is consistent with the fact that low-value import transactions are likely to be duty-free. [^14] Volpe-Martincus et al., (2024) get a similar result when comparing firms’ and consumers’ imports. The former would be equivalent to our high-value transactions and the latter to our low-value transactions. However, we observe that belonging to the EU has a negative effect on the share of low-value export transactions in the total number of export transactions, while it has a positive effect on imports.
@@
The fifth stylized fact, illustrated by panel B of Figure 2, is that the number of firms that only trade in low-value transactions is smaller than that that trade in high-value transactions, or in high and low-value transactions. For example, in 2023, the number of firms that only traded in low-value transactions represented 8% and 16% of all exporters and importers, respectively. These percentages were much higher than those at the beginning of the period (2% for both exporters and importers). There is an increase in the percentage of exporters that trade in large- and low-value transactions. In contrast, the percentage of this category of firms decreases among importers. We find a positive correlation between the number of large- and low-value transactions performed by a firm both for exports and imports. We also find a positive correlation between the total value of large- and low-value transactions (Table A.2 in the Appendix).
@@
Finally, the sixth stylized fact, illustrated by panel C of Figure 2, is that the concentration of the number of transactions by firm is higher in low-value transactions than in high-value ones. For example, the top firm in number of low-value export transactions represented 20% of all low-value export transactions in 2023, while the top firm in high-value export transactions “only” represented 6% of the total high-value export transactions. For imports, the differences were even larger: the top firm represented 21% and 2% of the total number of small- and high-value transactions, respectively. Furthermore, the concentration of transactions by firm increased between 1997 and 2023, in particular for low-value transactions: the share of the top firm in the total number of transactions increased from 3% to 20% in exports and from 3% to 21% in imports.